In a triangle, length of 2 larger sides is 24 & 22 resp.. if angle r in AP, then 3rd side is.........................


Also give reason behind choice of sides...........

Grade:12

3 Answers

Chetan Mandayam Nayakar
13 years ago

 

Dear Piyush, 

the angle opposite 22 is 60 degrees. Let third side be x,

from cosine law, 222=x2+242 -2*24*x*cos(60)

x2-24x+92=0

x=12±√52=12±2√13

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar respectively : Click here to download the toolbar.. 

Askiitians Expert

Chetan Mandayam Nayakar – IIT Madras  

vikas askiitian expert
13 years ago

let ABC be the triangle...

angles are in ap so 2B=A+C...........1

                          A+C=90-B..........2

         from 1 and 2

B=60 and so A and C can be 30 and 90

 now draw a triangle ABC right angled at C

    AB is the hypotenuse therefore AB=24

let the other sides be 22 and x

 now 24^2 =22^2 +x^2

         X=+-2sqrt23

Surya Anuraag Duvvuri
13 years ago

te angle opposite to 22 is 60 degrees,let the third side be a

now from cos rule 22^2=x^2+24^2-2*x*24 Cos60

                 x^2-24x+92=0

       x = 12+-rt52=12+-2rt13